### v3

# EMOE: A Framework for Out-of-distribution Uncertainty Based Rejection via Model-Agnostic Expansive Matching of Experts

Abstract
Expansive Matching of Experts (EMOE) is a novel framework that utilizes support-expanding, extrapolatory pseudo-labeling to improve prediction and uncertainty based rejection on out-of-distribution(OOD) points. EMOE utilizes a diverse set of multiple base experts as pseudo-labelers on the augmented data to improve OOD performance through multiple MLP heads (one per expert) with shared embedding train with a novel per-head matching loss. Unlike prior methods that rely on modality-specific augmentations or assume access to OOD data, EMOE introduces extrapolatory pseudo-labeling on latent-space augmentations, enabling robust OOD generalization with any real-valued vector data. In contrast to prior modality agnostic methods with neural backbones, EMOE is model-agnostic, working effectively with methods from simple tree-based models to complex OOD generalization models. We demonstrate that EMOE achieves superior performance compared to state-of-the-art method on diverse datasets in single-source domain generalization setting.
1 Introduction
It is well-known that the generalization capabilities of models can be severely limited when tested on out-of-distribution (OOD) data that deviates from the training-time distribution (Torralba and Efros 2011; Liu et al. 2021; Freiesleben and Grote 2023). This, in turn, affects many real-world applications where models may be evaluated on distribution-shifted data during deployment. For instance, these issues commonly arise in medical applications where patient distributions at inference time may deviate from the training data (Lee, Yin, and Zhang 2023). A potential strategy for the safe deployment of models in real-world applications is to employ novelty-based rejection (Dubuisson and Masson 1993; Hendrickx et al. 2024), where predictions are rejected whenever the model is evaluated on an instance that deviates from the data distribution seen during training. While such approach is appropriate in certain scenarios (e.g., when a human can easily intervene upon rejection), this prevalent strategy is overly conservative as it foregoes any potential extrapolation [^1] We use the term extrapolation to loosely encompass prediction outside of the training data distribution support (without consideration of the data’s convex-hull). by design. That is, novelty-rejection forbids any form of extrapolation (predictions outside of the training data support), even when the model may be capable.
Motivation & Broader Impacts   As a motivating example, consider a virtual screening application in drug discovery that aims to identify potential targets in molecules with machine learning, where a model must predict the properties of novel molecules that are often substantially different from previously characterized compounds. This requires models to generalize beyond the support of the training data, an ability referred to as ”scaffold hopping” (Hu, Stumpfe, and Bajorath 2017) to identify structually distinct candidates with desirable properties. However, most existing novelty-based rejection methods are designed to forbid extrapolation, restricting predictions to regions well-supported by the training data (Dubuisson and Masson 1993). This conservative approach, while ensuring reliability within known domains, fundamentally limits the discovery of novel candidates, which is the primary objective in drug discovery. Moreover, because experimental validation of candidate molecules is expensive and resource-intensive (Serrano et al. 2024), it is crucial to predict effectively on OOD instances while providing accurate predicted confidences. To manage cost only a small fraction of candidates should be selected for further laboratory tests. The use of models with poor confidence filtration will result in false positives that waste resources on unsuccessful experiments where screened molecules do not present desirable properties. Therefore, the success of discovery relies on the quality of high-confidence predictions on OOD data. We focus on the single source generalization setting (Qiao, Zhao, and Peng 2020) because it realistically reflects the constraints in practical virtual screening workflows where only a single labeled data set, often centered on a narrow chemical space, is available for training. Therefore, developing models that generalize from a single source to OOD drug candidates, while providing reliable confidence estimates, is a necessary direction for robust and cost-effective drug discovery.
[…]
Figure 1: Tally of datasets where respective methods lead in metrics: AUPRC for recall <.2<.2 (AUPRC@R<.2<.2), AUPRC, & AUROC. (See further details in $\S$4 and Appx.B.)
Figure 2: EMOE takes a 2-stage training approach. Stage 1 trains diverse base experts based on different subset of the training data (left). Stage 2 trains Multi-Headed Neural Network with pseudo labels on augmented data provided by the base experts (right).
Contributions  In this work, we propose a new method for single source domain generalization – EMOE, designed to extrapolate effectively and yield reliable uncertainty estimates in high-confidence regions through a novel training scheme using a multi-headed network based on matching augmented (expanded) data (see Fig. 2).
[…]
$\mathrm{AUPRC@R{<}\tau}=\frac{1}{\tau}\int_{0}^{\tau}\text{Precision}(r)\,dr$ (1)
This novel metric measures a model’s ability to predict confidence of true positive examples in its most confident predictions. (4) We show state-of-the-art (SOTA) performance in rejecting predictions based on estimated confidences, evaluated using AU{PR,RO}C-based metrics (see Fig. 1) in a single-source generalization setting (Qiao, Zhao, and Peng 2020). (5) We conduct several ablations to better understand the keys to EMOE’s success; moreover, by ablating the type of base experts, we show EMOE’s broad ability to improve over a model-agnostic set of experts.
2 Related Work
Domain Generalization  Domain generalization (DG) aims to learn a model that is able to generalize to multiple domains. A typical approach is to learn a domain invariant representation across multiple source domains. Domain invariant representation learning can be done by minimizing variations in feature distributions (Li et al. 2018; Ding et al. 2022) and imposing a regularizer to balance between predictive power and invariance (Arjovsky et al. 2019; Koyama and Yamaguchi 2020). Another line of research incorporates data augmentation to improve generalizability. Basic transformations like rotation and translation, varying in magnitude, are commonly used on images to diversify the training data (Cubuk et al. 2019; Berthelot et al. 2020). More sophisticated augmentation techniques have recently surfaced: (Zhang et al. 2018) introduced mixup, which linearly combines two training samples; (Yun et al. 2019) proposed CutMix, blending two images by replacing a cutout patch with a patch from another image; (Zhong et al. 2022) adversarially augment images to prevent overfitting to source domains. We focus on augmentations that are general and applicable across modalities. (Tian et al. 2022) introduced NCDG, which uses simple augmentations along with a loss function that maximizes neuron activity during training while minimizing standard classification loss. Their method minimizes the difference in the gradient of a coverage loss between standard training instances and augmented training instances.
Self-Training  Self-training uses an earlier model to pseudo-label unlabeled data, which is then added to the training set for subsequent model updates.The concept of pseudo-labeling was initially proposed by (Lee 2013), suggesting a direct approach to retaining instances where the model has high prediction probabilities. Following (Lee 2013), (Zou et al. 2018) proposed selecting a proportion of the most confident unlabeled points instead of using a fixed threshold. Later works combined pseudo-labeling with curriculum learning, dynamically adjusting class-wise thresholds over time to incorporate more informative samples (Cascante-Bonilla et al. 2020; Zhang et al. 2021).Another line of work improves pseudo-labeling robustness by promoting diversity in the labelers. (Ghosh et al. 2021) used model ensembles as teachers, while (Xie et al. 2019) added noise via Dropout (Srivastava et al. 2014) and data augmentation. FixMatch (Sohn et al. 2020) generates pseudo-labels from weakly augmented samples to supervise training on the corresponding strongly augmented samples. EMOE novelly leverages pseudo-labels by assigning each student head to a different expert, encouraging diversity, while aligns the ensemble prediction for consistency via a novel loss.
Selective Classification  Reject option methods (or selective classification) aim to identify inputs where the model should abstain from predicting. Many approaches apply post hoc processing: after training, a rejection metric—such as the model’s predicted probability—is computed, and predictions below a set threshold are rejected (Stefano, Sansone, and Vento 2000; Fumera, Roli, and Giacinto 2000). Building upon these works, (Devries and Taylor 2018) proposed to train a confidence branch alongside the prediction branch by incentivizing a neural network to produce a confidence measure during training; (Geifman and El-Yaniv 2017) proposed a method for constructing a probability-calibrated selective classifier with guaranteed control over the true risk. Recently, methods adopting end-to-end training approaches have been proposed (Thulasidasan et al. 2019; Ziyin et al. 2019; Geifman and El-Yaniv 2019). In these works, an extra class is added when predictions are made. If the extra class has the highest class probability for a sample, the sample is rejected. Most reject-option approaches are geared towards in-distribution rejection and utilize novelty-rejection when encountering any OOD points (Torralba and Efros 2011; Liu et al. 2021; Freiesleben and Grote 2023); instead, we propose to learn better conditional output probabilities on OOD data for more effective, capability-aware rejection.
Ensemble Modeling  Ensembles utilize a diverse set of models jointly for better performance. Early methodologies for ensembles aggregate (bag) predictions from all models (Dietterich 2007; Kussul et al. 2010) or a subset of the models in the ensemble (Jordan and Jacobs 1993; Eigen, Ranzato, and Sutskever 2013). In the OOD setting, prior works addressed this problem by enforcing prediction diversity on OOD data (Pagliardini et al. 2023), ensembling moving average models (Arpit et al. 2022a), training an ensemble of domain specific classifiers (Yao et al. 2023), and training diverse model heads within a single network by maximizing disagreement on unlabeled OOD data (Lee, Yao, and Finn 2022). EMOE adopts a multi-headed architecture that produces an ensemble to improve predictions on OOD data.
3 Method
Our approach, EMOE, consists of three core components: (1) training a diverse set of base experts; (2) generating extrapolatory samples via latent space augmentation; and (3) training a multi-headed network to match expert predictions on both ID and expanded data. This section describes each step in detail. Throughout, we assume the ‘single-source’ generalization setting (Qiao, Zhao, and Peng 2020), where we observe a single in-distribution (ID) training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and instances are drawn iid $(x_{i},y_{i})\sim\mathcal{P}_{\mathrm{in}}$ without any accompanying environmental/domain/source information nor any labeled/unlabeled OOD instances. For simplicity, we write to the binary classification case, $y_{i}\in\{0,1\}$, but our methodology is easily extendable to other supervised tasks. We design our method to work in general, non-modality specific [^2] We avoid any modality or domain-specific augmentations. (e.g., image, text, audio) settings, i.e., $x_{i}\in\mathbb{R}^{d}$.
3.1 Base Collection of Experts
EMOE leverages a set of diverse initial experts $\{g_{k}\}_{k=1}^{K}$, s.t. $g_{k}:\mathbb{R}^{d}\rightarrow\{0,1\}$, to guide the training of a secondary model by providing pseudo-labels. There are many mixture of experts (Jordan and Jacobs 1993; Eigen, Ranzato, and Sutskever 2013) and ensembling (Arpit et al. 2022a; Dietterich 2007; Pagliardini et al. 2023; Yao et al. 2023) methods available , EMOE utilizes a collection of strong base-learners by sub-selecting on both instance and feature subspace, encouraging diversity by specializing experts on distinct regions and views of the latent representation space.
3.2 Expansive Augmentation of Training Data
Figure 3: Expansion in UMAP space: ID points (orange) are augmented (purple) and expand the distributional support towards the OOD (green).
To train models capable of extrapolating to OOD samples, we need to expose them to data that lie outside the support of the training distribution. To reason about the support of the training data, and how to expand past it, we propose to leverage a latent factor space, $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}^{s}$. While learning semantically meaningful latent factor spaces remains an active area of research, we observed strong performance utilizing autoencoding techniques (see $\S$ 4), which carry a corresponding decoder $\gamma:\mathbb{R}^{s}\mapsto\mathbb{R}^{d}$. Without loss of generality, we consider centered latent spaces such that $\mathbb{E}[\varphi(X)]=0$.
We propose a novel, yet straightforward strategy to expand data outside of training distributional support: perturb instances to lie further away from the origin in latent space. In particular, if we have latent vector $z=\varphi(x)$, we propose to consider perturbations of the form $z^{\prime}=(1+|\epsilon|)z$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, and one can utilize the decoder $x^{\prime}=\gamma(z^{\prime})$. I.e,, we define our expansion operation on a set of points as:
$\textbf{Ex}(\{x_{i}\}_{i=1}^{N})\equiv\left\{\gamma\left((1+|\epsilon_{i}|)\varphi(x_{i})\right)\mid\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})\right\}_{i=1}^{N}.$ (2)
Ex will be a stochastic mapping. As shown in Fig. 3, our expansion covers areas away from the training support, even covering areas of OOD data. However, unlike with small jitter-based perturbations, where one can retain an original instance label, it is less clear how to derive an accompanying training signal for expansive augmentations. Below, we propose to leverage a pseudo-labeling scheme where we derive $K$ labels with the base experts $(f_{1}(x^{\prime}),\ldots,f_{K}(x^{\prime}))$ [^3] One may also train base experts directly on the latent space, $(f_{1}(z^{\prime}),\ldots,f_{K}(z^{\prime}))$, and avoid the decoder..
3.3 EMOE: Expansive Matching of Experts
Once we generate extrapolated data via latent expansion, the key challege becomes how to provide supervision on these OOD samples. Since true labels are unavalable, we leverage predictions from base experts to pseudo label these points. To retain the diversity in the base experts, we propose a multi-headed neural architecture trained via self-training. Specifically, the network consists of a shared multilayer perceptron (MLP), denoted as $\phi:\mathbb{R}^{d}\to\mathbb{R}^{m}$, which learns a common representation space, and $K$ expert-specific heads, $h_{1},\ldots,h_{K}$, where each head maps the shared representation to a prediction, e.g., $h_{j}:\mathbb{R}^{m}\to\mathbb{R}$ for binary classification. Importantly, each head is trained to match the output of a different base expert, enabling the model to align with multiple diverse supervisory signals. The training loss is defined as a sum of per-head losses on a pseudo-labeled set $\mathcal{S}$, where each $h_{j}$ is optimized to match the corresponding expert’s predictions. This encourages shared representation to support generalization across diverse pseudo-labeling sources. Our per head loss that matches pseudo labels provided by the experts to the MLP heads on a set $\mathcal{S}$ is as follow:
$\mathcal{L}_{\mathrm{match}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{S})$ (3)
$\equiv\frac{1}{|\mathcal{S}|K}\sum_{x\in S}\sum_{j=1}^{K}\ell(h_{j}(\phi(x)),g_{j}(x)),$ (4)
where $\ell(\hat{y},y)$ is a supervised loss (e.g., the cross-entropy loss). Moreover, we will utilize a mean-matching L1 loss
$\mathcal{L}_{\mathrm{mean}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{S})$ (5)
$\equiv\frac{1}{|\mathcal{S}|}\sum_{x\in S}\left\|\frac{1}{K}\sum_{j=1}^{K}\sigma(h_{j}(\phi(x)))-\frac{1}{K}\sum_{j=1}^{K}g_{j}(x))\right\|_{1},$ (6)
where $\sigma(\cdot)$ is the sigmoid. Our full expansive matching of experts’ loss is then:
$\mathcal{L}_{\mathrm{EMOE}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{D})$ (7)
$\equiv$ $\ \mathcal{L}_{\mathrm{mean}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{D})$ (8)
$+$ $\mathcal{L}_{\mathrm{match}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{D})$ (9)
$+$ $\lambda\,\mathcal{L}_{\mathrm{match}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\textbf{Ex}(\mathcal{D})).$ (10)
Note that we provide additional supervisory losses on non-augmented $\mathcal{D}$ via $\mathcal{L}_{\mathrm{mean}}$. Empirical results show ($\S$ 4) that the network heads learn often learn a better estimator than the base experts. However, we see more consistent improvements by not discarding the base experts and bagging:
$f_{\mathrm{EMOE}}(x)\equiv\frac{1}{2K}\sum_{j=1}^{K}g_{j}(x)+h_{j}(\phi(x)).$ (11)
Motivation
We expound on how EMOE may learn better estimates on OOD data through diversity and multi-task learning, and variance reduction and regularization.
Diversity and Multi-task Learning. In practice, we propose simple linear heads. At an intuitive level, this forces the MLP to learn a robust feature embedding that can ‘mimic’ the diverse views that the base experts provide. That is, this will force the last hidden layer to featurize an embedding $\phi(x)$ that can, with simple linear projections, emulate a diverse set of labels. The per-head matching loss (4) may be formulated as a multi-task loss on a set a of $K$ virtual environments $\mathcal{E}_{j}(\mathcal{S})=\{(x,g_{j}(x))\mid x\in\mathcal{S}\}$: $\mathcal{L}_{\mathrm{match}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{S})=\frac{1}{K}\sum_{j=1}^{K}\mathcal{L}(h_{j}(\phi(\cdot)),\mathcal{E}_{j}(\mathcal{S})),$ where $\mathcal{L}(h_{j}(\phi(\cdot)),\mathcal{E}_{j}(\mathcal{S}))$ is the supervised loss on instances/labels in environment $\mathcal{E}_{j}(\mathcal{S})$ with estimator $h_{j}(\phi(\cdot))$ on the shared embedding $\phi$. Thus, when training on the expanded set of data-points, $\textbf{Ex}(\mathcal{D})$, with pseudo-labels stemming from diverse experts (e.g., trained on different subsets of features and instances), we see that our matching loss provides supervisory signals to learn: 1) on OOD data (through expansion); 2) robust embeddings that must generalize to diverse environments.
Variance Reduction and Regularization. Previous work has decomposed OOD generalization into bias/variance terms (Yang et al. 2020; Arpit et al. 2022b):
$\mathbb{E}_{(x,y)\sim\mathcal{P}_{\mathrm{out}}}\mathbb{E}_{\mathcal{D}\sim\mathcal{P}_{\mathrm{in}}}[\mathrm{CE}(y,f(x;\mathcal{D}))]$ (12)
[…]
where $\mathrm{CE}$ is the cross-entropy loss, $f(x;\mathcal{T})$ is the model fit on dataset $\mathcal{T}$, $\bar{f}(x)=\mathbb{E}_{\mathcal{D}}[f(x;\mathcal{D})]$ is the expected prediction when averaging out draws on the (in-distribution) training dataset $\mathcal{D}$, and $\mathcal{P}_{\mathrm{out}}$ is the OOD data distribution at inference time. Letting $\bar{g}(x)\equiv\frac{1}{K}\sum_{j=1}^{K}g_{j}(x)$, we may view $\bar{g}(x)$ as a bootstrap-like estimate for $\bar{f}(x)$. One may then take $\mathbb{E}_{x,\mathcal{D}}[\mathrm{KL}(\bar{g}(x),f(x;\mathcal{D}))]$ as a proxy for $\mathbb{E}_{x,\mathcal{D}}[\mathrm{KL}(\bar{f}(x),f(x;\mathcal{D}))]$ and roughly consider
$\mathbb{E}_{(x,y)\sim\mathcal{P}_{\mathrm{out}}}\mathbb{E}_{\mathcal{D}\sim\mathcal{P}_{\mathrm{in}}}[\mathrm{CE}(y,f(x;\mathcal{D}))]$ (14)
[…]
which connects to (7) when interpreting our expanded points as a proxy for the (unknown) OOD distribution $\mathcal{P}_{\mathrm{out}}$ and $\mathcal{L}_{\mathrm{match}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\textbf{Ex}(\mathcal{E}))$ as a proxy for $\mathbb{E}_{x,\mathcal{D}}[\mathrm{KL}(\bar{g}(x),f(x;\mathcal{D}))]$.
4 Experiments
Table 1: Experiment results on DrugOOD (Ji et al. 2023) datasets. We bold best scores based on the mean minus 1 standard error. We italicize best scores when they are achieved by a semi-supervised method (*), that uses additional unlabeled OOD data during training.
refined ec50 val refined ec50 test core ic50 val core ic50 test core ec50 val core ec50 test
AUPRC@ DivDis 96.02±0.19 86.51±0.28 97.80±0.07 89.72±0.06 93.86±0.42 80.18±0.92
R<<0.2 FixMatch 94.48±0.47 86.38±0.35 97.62±0.06 92.89±1.36 95.02±0.76 70.49±0.43
D-BAT 96.97±0.16 88.78±0.40 98.13±0.08 91.79±0.38 93.81±0.22 84.35±1.35
AdvStyle 95.13±0.13 88.21±0.37 97.04±0.17 89.05±0.22 94.84±0.31 84.51±2.36
EoA 85.03±0.06 78.79±0.14 88.56±0.05 77.03±0.13 81.85±0.24 71.84±0.45
Mixup 85.39±0.23 79.78±0.34 88.99±0.43 78.07±0.61 83.97±0.61 73.04±0.42
NCDG 93.92±0.62 84.46±0.67 97.83±0.08 87.82±0.26 93.40±0.78 80.09±1.94
EMOE base 98.00±0.07 89.48±0.24 99.14±0.02 94.20±0.15 97.79±0.11 68.48±0.30
EMOE 98.45±0.06 89.76±0.26 99.15±0.05 94.42±0.09 98.66±0.10 69.04±0.50
AUPRC DivDis 89.62±0.12 80.92±0.06 93.34±0.10 81.48±0.30 83.85±0.41 75.09±0.38
FixMatch 87.14±0.58 81.51±0.31 92.38±0.15 82.78±0.57 80.92±1.12 70.57±0.32
D-BAT 84.70±0.51 70.08±0.69 90.84±0.28 73.45±0.90 76.64±0.49 54.87±0.99
AdvStyle 83.01±0.90 69.48±2.16 88.54±1.44 72.11±1.49 81.17±1.72 58.40±2.18
EoA 69.66±0.47 57.71±0.79 79.12±0.09 56.52±0.42 64.16±0.51 36.50±1.48
Mixup 80.36±0.88 72.88±1.67 86.88±0.14 74.99±0.15 73.03±1.67 60.84±4.31
NCDG 89.27±0.22 80.54±0.21 94.17±0.08 81.19±0.11 87.72±0.37 74.99±0.74
EMOE base 91.21±0.02 82.61±0.05 94.91±0.03 84.13±0.06 88.44±0.05 71.80±0.07
EMOE 91.59±0.03 83.06±0.10 95.38±0.01 84.77±0.02 89.52±0.05 71.41±0.11
We conduct experiments on a varied set of real-world datasets to test the OOD generalizability of EMOE. We considered the single source domain generalization setting (e.g., (Qiao, Zhao, and Peng 2020)), where our model is trained solely on ID data without any (labeled or unlabeled) OOD data during training/validation (e.g., precluding typical semi-supervised approaches), and without any accompanying environmental/domain/source information from ID training instances. Moreover, we note that we avoided utilizing any modality-specific information in EMOE (e.g., we do not utilize any domain specific augmentations) for generality. We utilized XGB Classifiers (Chen and Guestrin 2016) fitted to random subsets of data instances and features as the base collection of experts. For a fair/realistic evaluation, we avoided any hyper-parameter tuning on EMOE and utilized a fixed architecture of a $2$ layer $512$ ELU (Clevert, Unterthiner, and Hochreiter 2015) hidden-unit MLP with $1024$ linear-output heads (please see other hyperparameters in Appx.A.1). For our latent space, we utilize PCA with $128$ components. While OOD generalization is an active field of research (Freiesleben and Grote 2023; Liu et al. 2021), methodology for general (non-modality specific) single source domain generalization is more limited. We provide context to our results through comparisons to a diverse set of existing strong domain generalization methods that approach the problem from various perspectives (and are applicable in the single-source setting).
Single-Source Baselines  In our experiments, we include three baselines that utilize data augmentation: AdvStyle (Zhong et al. 2022), Mixup (Zhang et al. 2018), and NCDG (Tian et al. 2022) as well as two baselines employing ensemble methods: D-BAT (Pagliardini et al. 2023) and EoA (Arpit et al. 2022a). Mixup linearly combines two ID samples, AdvStyle adversarially augments ID data, NCDG takes a simple augmentation and optimizes neuron coverage, D-BAT enforces prediction diversity on OOD data, and EoA ensembles moving average models.
Semi-Supervised Baselines  We provide further context by comparing to semi-supervised methods for general tabular data, DivDis (Lee, Yao, and Finn 2022) and FixMatch (Sohn et al. 2020). Note, these methods utilize more information than EMOE, and have access to unlabeled OOD data from the test-time distribution. That is, these approaches break from our single-source generalization setting and utilize additional information of the test-time distribution. DivDis is a two-stage framework that first trains diverse model heads using unlabeled target OOD data and then selects the best head with minimal supervision (i.e., using a single label from target OOD data). FixMatch generates filtered pseudo-labels using the model’s predictions on weakly-augmented OOD unlabeled samples. The model is then trained to predict the pseudo-label when provided with a strongly augmented version of the same sample. As DivDis and FixMatch utilize additional information from single-source methods, we separate their results and denote them with an asterisk.
Figure 4: AUPRC at recall << 0.2 for EMOE and competitive baselines on hERG dataset.
Metrics  For prediction thresholding (rejection), we directly utilize the conditional probability $P(Y=1\mid X=x)$ generated by the models. Many real-world applications (e.g. drug virtual screening) utilize only high-confidence predictions. Thus, we paid close attention to high-confidence filtration and reported percent of AUPRC at conservative recall thresholds $\mathrm{AURPC@R}\!\!<\!\!\tau$’ (see Eq.1 for formal definition and Fig.4 for illustration). We also reported AUPRC, AUROC (see Appx. B.1 & B.2 for results on AUROC, which follow a similar trend) We report both base expert (‘EMOE Base’) and EMOE (Eq.11) ensemble performance. Our code will be open-sourced upon publication.
Table 2: Experiment results on ChEMBL (Gaulton et al. 2011) and Therapeutics Data Commons (Huang et al. 2021) datasets. We bold best scores based on the mean minus standard error.
hERG A549_cells cyp_2D6 Ames
AUPRC@ DivDis 85.65±2.37 89.45±1.16 79.98±1.10 95.80±0.34
R<<0.2 FixMatch 79.66±0.57 90.81±4.30 83.58±1.98 93.55±0.64
D-BAT 84.48±1.74 98.26±0.14 91.40±0.98 99.04±0.24
AdvStyle 88.21±0.77 97.77±0.28 84.83±0.93 99.05±0.17
EoA 63.80±0.42 61.31±0.31 61.77±0.41 78.74±0.43
Mixup 82.25±1.51 95.04±0.25 87.09±2.32 91.02±1.06
NCDG 78.25±2.60 90.33±0.95 79.31±2.44 87.86±1.41
EMOE base 94.44±0.17 98.22±0.08 95.51±0.19 97.84±0.20
EMOE 94.67±0.29 98.87±0.09 96.88±0.25 98.45±0.19
AUPRC DivDis 67.70±0.25 76.45±0.63 65.76±0.45 82.37±0.27
FixMatch 45.16±3.11 51.31±1.28 32.65±2.16 72.54±1.15
D-BAT 54.60±1.59 67.04±0.53 47.42±0.85 70.44±0.76
AdvStyle 51.54±0.93 65.02±0.72 44.41±0.96 74.98±0.49
EoA 43.30±0.51 44.95±0.24 37.37±0.95 59.43±0.18
Mixup 42.42±0.85 50.52±0.50 27.79±1.49 60.94±0.80
NCDG 65.03±0.75 79.50±0.79 65.31±1.00 76.48±0.30
EMOE base 72.19±0.07 84.10±0.01 72.93±0.13 87.43±0.02
EMOE 73.26±0.08 84.60±0.08 73.59±0.15 88.50±0.10
4.1 Chemical Datasets
To test how well EMOE generalizes to OOD data in chemical domains, we considered seven datasets from ChEMBL (Gaulton et al. 2011), Therapeutics Data Commons (Huang et al. 2021), and DrugOOD (Ji et al. 2023). For all datasets, we represented molecules using extended-connectivity fingerprints (Rogers and Hahn 2010) with radius 2 (ECFP4) and with dimensionality 1024. ECFP4 is a standard method for molecular representation and was chosen for its simplicity in calculation as well as its ability to perform comparably to learned representations, such as those generated by graph neural networks on relevant classification tasks (Zagidullin et al. 2021). Datasets for inhibition of human Ether-à-go-go-Related Gene (hERG), cytotoxicity of human A549 cells (A549_cells), and agonists for Cytochrome P450 2D6 (cyp_2D6) were collected from ChEMBL (Gaulton et al. 2011). For these datasets, binary classification labels were generated with a pChEMBL threshold of 5.0. We also considered another binary classification dataset for Ames mutagenicity (Ames) that was taken from Therapeutics Data Commons (TDC) (Huang et al. 2021). For the ChEMBL and TDC datasets (hERG, A549_cells, cyp_2D6, and Ames), ID and OOD splits were determined based on the Murko scaffold of a molecule, such that OOD data have molecular scaffolds not present in the ID data, mimicking the “scaffold domain” approach utilized in DrugOOD (Ji et al. 2023).
Moreover, we considered the “core ec50,” “refined ec50,” and “core ic50” ligand-based affinity prediction datasets (lbap) from DrugOOD (Ji et al. 2023) (the three hardest OOD performance gap datasets). For these datasets ID and OOD splits were determined based on the number of atoms in a molecule, such that larger molecules are considered the OOD set and smaller molecules, the ID set. Datasets are organized by domain and subsequently divided into training, OOD validation, and OOD testing sets in sequential order. While larger in samples, the DrugOOD datasets ignore the impact of biological targets on label, resulting in a modeling task that has limited relevance to drug discovery. The additional ChEMBL/TDC datasets, though smaller, have direct relevance for drug discovery.
Figure 5: Predicted probabilities from EMOE network and base experts. We highlight example instances where the base experts initially makes incorrect predictions but are corrected when we average the predicted probabilities from EMOE network and EMOE base.
Results  We assessed models on ChEMBL, TDC, and DrugOOD datasets; our results are shown in Tab. 2 and Tab. 1. Across the chemical datasets, EMOE consistently outperforms both supervised and semi-supervised baselines and consistently shows gains over its base experts. Note further, that EMOE’s performance gain is especially impactful at conservative recalls (AUPRC@R<<0.2, e.g., see Fig. 4), which indicates the EMOE outperform other models in virtual screening tasks to filter drug candidates (where false positives would lead to wasted resources). Please see discussion below ($\S$ 4.4), for further analysis on the improved performance of EMOE w.r.t. increased confidence (Fig. 5) and expert diversity (Fig. 6).
4.2 Other Real World Datasets
Table 3: Experiment results on Tableshift (Gardner, Popovic, and Schmidt 2023) datasets. We bold best scores based on the mean minus 1 standard error.
Childhood FICO Hospital Sepsis
Lead HELOC Readmission
AUPRC@ DivDis 89.77±2.51 85.50±2.65 83.22±2.57 60.66±1.57
R<<0.2 FixMatch 81.75±3.69 77.76±1.94 67.84±1.95 17.28±0.58
D-BAT 62.82±0.00 91.20±0.24 78.84±0.12 75.37±0.38
AdvStyle 64.96±0.01 88.71±0.65 72.91±0.58 59.83±0.63
EoA 77.43±0.97 59.53±2.24 51.83±1.66 41.10±1.56
Mixup 50.00±0.00 91.16±2.10 69.19±3.95 66.09±1.15
NCDG 42.49±1.31 87.93±2.35 68.88±1.15 15.22±0.38
EMOE base 97.39±0.06 90.07±0.32 58.30±0.09 78.62±1.57
EMOE 97.92±0.20 91.72±0.62 67.57±0.12 76.95±1.45
AUPRC DivDis 76.33±0.86 82.47±1.05 65.95±2.15 58.85±1.21
FixMatch 75.39±0.58 72.22±4.46 56.54±1.21 18.24±0.66
D-BAT 71.85±0.01 80.91±0.17 63.29±0.08 58.17±0.21
AdvStyle 48.37±2.77 79.63±1.65 38.13±3.80 54.34±0.27
EoA 49.48±0.19 59.53±2.24 29.45±4.95 11.21±2.21
Mixup 50.00±0.00 80.95±0.63 14.15±1.06 56.80±0.40
NCDG 23.08±0.24 79.05±1.21 58.95±0.23 11.96±0.32
EMOE base 86.39±0.19 83.80±0.07 62.83±0.03 64.01±0.94
EMOE 86.70±0.28 84.02±0.09 63.62±0.08 62.21±0.52
Next, we further evaluate our method in non-chemical domains across a diverse range of real-world OOD scenarios using the Tableshift datasets (Gardner, Popovic, and Schmidt 2023) (and the PACS image dataset (Li et al. 2017), see  Mat. B.5). We selected a diverse collection of Tableshift datasets, based on unrestricted availability and in/out-of-domain performance discrepancy, coverings areas including: finance, education, and healthcare. Each dataset has an associated real-world shift and a related prediction target (see Gardner, Popovic, and Schmidt (2023) for further details). Results on the Tableshift are shown in Tab. 8. As before, we consider the same single-source domain generalization setting. We can see that even over diverse applications, our EMOE method is able to perform well and is often outperforming our strong competing baselines. Moreover, eventhough DivDis and FixMatch use additional unlabeled OOD data, EMOE outperforms both on the majority of datasets, highlighting EMOE’s ability for generalization on completely unforeseen OOD instances (without any knowledge of test time distribution).
4.3 Ablation Studies
Model Agnostic Property
Table 4: ChEMBL Datasets’ Mean @R<0.2 Ablating Type of Base Experts.
XGBoost Random Forest Decision Tree D-BAT
Base 95.87 95.72 92.18 91.63
EMOE 96.64 96.15 94.15 93.81
Despite its dependency on base experts for pseudo-label generation, EMOE consistently improves performance across a wide range of base model types, including XGBoost, Random Forest, Decision Tree, and the OOD generalization method D-BAT (Pagliardini et al. 2023). As shown in Tab.4 (see details on this ablation in Appx.B.6). EMOE consistently improves upon each kind of base model, demonstrating its effectiveness as a model-agnostic framework. Moreover, this shows the future promise of incorporating EMOE with complementary methodology for OOD generalizability.
Per-head Matching Ablation Next, we ablate on the per-expert matching loss scheme (eq.7) used in EMOE with a simpler alternative: mean-only matching (MM) (eq.8) on the expanded points. This ablation tests whether the diversity induced by per-head supervision aids performance. Here we utilized single headed (SH) and multi-headed (MH) MLPs with mean matching, where EMOE uses multi-headed MLP with per expert loss (see  Mat.B.7 for details on the alternative losses). EMOE achieves the highest AUPRC improvements ($\Delta$) over base experts (+1.11 at R<<.2, +1.03 at R<<1) compared to SH+MM (+0.59, +0.45) and MH+MM (+0.25, +0.55), showing evidence for the effectiveness of training our embedding through the diverse multi-task loss induced by per-head matching ($\S$ 3).
Bottleneck Ablation  We motivated EMOE’s performance in terms of a multi-task scheme, which bottlenecks (the last hidden layer) to learn an embedding that can generalize to mimic predictions on a diverse set of pseudo-labels ($\S$ 3). We test this motivation by comparing the original architecture (Full: $2\times 512$ hidden layer) with a stronger bottleneck still, a small architecture (Tiny: $2\times 32$ hidden layer). (See Appx. B.8 for more details and results.) The performance gap is marginal between the ‘Full’ and ‘Tiny’ model (a $0.85\%$ difference) when using our proposed loss. In contrast, when using empirical risk minimization, we see a $4.87$ times bigger drop in performance between ‘Full’ and ‘Tiny’ models. This suggests that the bottlenecking properties of our method are key to EMOE’s performance, and show promise for EMOE in resource-constrained settings.
4.4 Discussion
Figure 6: Experts correlations between EMOE base experts and EMOE network heads on “Ames” dataset. (a) EMOE experts have a high correlation with EMOE base experts on sample the base experts make correct predictions. (b) EMOE experts show low correlations with EMOE base experts on samples where the base experts make incorrect predictions.
EMOE addresses the challenging single source OOD generalization problem for real-valued vector data, outperforming baselines, including those that leverage unlabelled OOD data. Across all metrics/datasets, EMOE achieves leading tally of 29, far exceeding both EMOE base and the next highest score of 4 from the strong competing baseline D-Bat (Pagliardini et al. 2023) as shown in Fig. 1. Despite its performance, EMOE does not incur an out-sized computational cost; the training of base experts may be done in parallel, and the cost of expanding data in the latent space is negligible (see Appx. A.3 for addition timing details).
In summary, EMOE presents a simple yet powerful model-agnostic approach to OOD generalization under single-source setting via pseudo-labels and expert-specific matching. EMOE’s architecture promotes a robust embedding and stands out in high confidence areas for OOD predictions. We believe this work opens promising directions for general-purpose, modality- and model- agnostic OOD learning, particularly in high-stakes applications like drug discovery, where confident extrapolation is critical and labeled data is scarce.
Appendix A Addtional Experiment Details
A.1 EMOE training details
In all of our experiments we used the Adam (Kingma and Ba 2014) optimizer and mini-batches of size 256. One Nvidia A100 GPU with 40GB GPU memory was used to run our experiments, and duration for model training is approximately 0.5 hours. $\lambda$ = 0.5 was used for the $\mathcal{L}_{\mathrm{match}}$ for the expanded points. As noted in Sec. 3.2 we trained the EMOE models directly in the latent space to avoid the need for the decoder (and also allowed baselines to do this if it aided their performance). In the experiments on hERG, A549_cells, CYP_2D6, Ames, core ec50, refined ec50, EMOE was trained for 20000 iterations. Arithmetic mean between EMOE and EMOE base was reported. We performed 5 trails on each of the datasets for EMOE.
A.2 Baseline Setup
We implemented all baselines we are comparing against EMOE following the implementation details in their paper and/or using Github implementations (if available). Since the fingerprints representation of chemicals are quite sparse, we preformed dimension reduction using PCA with 128 components on all chemical datasets. For D-BAT (Pagliardini et al. 2023) with existing implementations designed for tabular data, we utilized their original model architectures. For the other three baseline methods without implementation specifically for tabular data, we adopted a structure comprising two 512 ELU(Clevert, Unterthiner, and Hochreiter 2015) layers to closely mimic the EMOE network architecture. The Adam (Kingma and Ba 2014) optimizer was used for training baseline models.
D-BAT In our experiments, the D-Bat(Pagliardini et al. 2023) models used MLP architecture with one 128 LeakyRelu(Maas 2013) layer following the architecture in their Github. Their paper (Pagliardini et al. 2023) discussed two settings, and we focused on the scenario where perturbation data differs from the distribution of test data, adhering to the single-source domain generalization setting. We trained an ensemble of five models sequentially for the D-bat baseline models and the predictions from the 5 models were averaged to obtain the final prediction.
EoA We trained an ensemble of 5 simple moving average model following the method described in (Arpit et al. 2022a). We start calculating the moving average at iteration 50 and trained the models for 200 iterations. The predictions from the 5 models were averaged to obtain the final prediction for EoA.
For AdvStyle (Zhong et al. 2022) and Mixup (Zhang et al. 2018), the methodologies were straightforward. We experimented with training using various numbers of iterations and reported the most promising results. Note that we used alpha=0.7 when combining the 2 samples for Mixup. We executed all baseline experiments five times on each dataset to ensure a precise estimation of performance.
DivDis We utilized all unlabeled target OOD data for training and a single label from this data for supervision. An ensemble of 5 models was trained for 100 iterations with early stopping, and each model has 2 classification heads. Across all datasets, we set $\lambda_{1}=10$ (encouraging disagreement among model heads), while $\lambda_{2}$ (an optional hyperparameter prevents degenerate solutions) was set to 0 for DrugOOD and ChEMBL and to 10 for TableShift. The final prediction is the average of the 5 models’ predictions.
FixMatch Originally, Fixmatch (Sohn et al. 2020) was designed for image data, so the sense of weak and strong augmentations were image based. To adapt the method to our modality agnostic setting we used $x*(1+\alpha)$ as the weak augmentation, $x*(1+2*\alpha)$ as strong augmentation, where $\alpha$ is a small noise drawn from the standard normal distribution.
For NCDG (Tian et al. 2022), we adapt the method to use the EMOE architecture (rather than a ResNet model) and the EMOE augmentation method. We set t=0.005 (the threshold for neuron activation in coverage computation), $\lambda=0.1$ (the weight coefficient for neuron coverage loss), and $\beta=0.01$ (the weight for gradient similarity regularization loss). Five trials were run on each dataset and averaged to obtain the final results.
A.3 EMOE training time
In Tab. 5, we report the training time for EMOE and baseline models we are considering. Note that the base experts can be training in parallel with enough computational resources (and are each quick to train at <1<1s).
Table 5: Training Time for EMOE and baseline methods.
wall clock (m)
EMOE NN 4.2
Dbat 3.5
EoA 67.9
Advstyle 4.8
Mixup 1.2
NCDG 2.7
Appendix B Additional Experiment and Ablation Results
Table 6: Full experiment results on DrugOOD datasets. We bold best scores based on the mean minus 1 standard deviation. Note that * refers to a semi-supervised method.
core ec50 val core ec50 test refined ec50 val refined ec50 test core ic50 test core ic 50 test
AUPRC@ DivDis 96.21±1.02 82.45±2.93 96.63±0.54 88.87±0.58 98.41±0.26 91.72±1.47
R<<0.1 FixMatch 96.42±1.06 69.80±0.25 94.78±0.63 86.53±0.25 97.70±0.05 95.89±1.46
D-BAT 94.04±0.25 86.59±1.42 97.19±0.19 88.93±0.48 98.25±0.09 91.89±0.39
AdvStyle 95.79±0.20 84.56±2.36 96.37±0.32 88.69±0.43 98.10±0.17 89.39±0.33
EoA 81.85±0.24 71.84±0.45 85.03±0.06 78.79±0.14 88.56±0.05 77.03±0.13
Mixup 83.97±0.61 73.03±0.42 85.39±0.23 79.78±0.34 88.99±0.43 78.07±0.61
NCDG 91.95±0.28 80.38±0.47 96.33±0.07 89.94±0.12 97.36±0.08 89.68±0.05
EMOE base 98.73±0.14 65.40±0.52 99.00±0.04 92.67±0.33 99.57±0.01 96.93±0.09
EMOE 99.06±0.14 64.71±0.29 99.22±0.05 91.31±0.41 99.36±0.06 96.42±0.10
AUPRC@ DivDis 93.86±0.42 80.18±0.92 96.02±0.19 86.51±0.28 97.80±0.07 89.72±0.06
R<<0.2 FixMatch 95.02±0.76 70.49±0.43 94.48±0.47 86.38±0.35 97.62±0.06 92.89±1.36
D-BAT 93.81±0.22 84.35±1.35 96.97±0.16 88.78±0.40 98.13±0.08 91.79±0.38
AdvStyle 94.84±0.31 84.51±2.36 95.13±0.13 88.21±0.37 97.04±0.17 89.05±0.22
EoA 81.85±0.24 71.84±0.45 85.03±0.06 78.79±0.14 88.56±0.05 77.03±0.13
Mixup 83.97±0.61 73.04±0.42 85.39±0.23 79.78±0.34 88.99±0.43 78.07±0.61
NCDG 89.82±0.22 77.94±0.39 95.87±0.05 88.60±0.10 96.42±0.06 87.89±0.04
EMOE base 97.79±0.11 68.48±0.30 98.00±0.07 89.48±0.24 99.14±0.02 92.31±0.14
EMOE 98.66±0.10 69.04±0.50 98.45±0.06 89.76±0.26 99.15±0.05 92.95±0.05
AUPRC@ DivDis 91.89±0.95 78.97±1.66 95.19±0.37 85.12±0.49 97.33±0.16 88.06±1.14
R<<0.3 FixMatch 93.76±1.11 70.61±0.47 87.14±0.64 85.74±0.42 97.59±0.07 91.02±1.24
D-BAT 93.73±0.21 81.40±0.89 96.89±0.15 87.77±0.32 98.08±0.08 90.71±0.52
AdvStyle 94.52±0.37 83.57±2.08 94.71±0.13 88.05±0.36 96.69±0.21 88.93±0.19
EoA 81.85±0.24 71.84±0.45 85.03±0.06 78.79±0.14 88.56±0.05 77.03±0.13
Mixup 83.97±0.61 73.06±0.43 85.39±0.23 79.78±0.34 88.99±0.43 78.07±0.61
NCDG 88.73±0.19 76.05±0.35 95.17±0.04 86.07±0.08 95.89±0.05 86.53±0.03
EMOE base 96.73±0.10 69.94±0.20 97.17±0.06 87.99±0.18 98.71±0.03 92.32±0.23
EMOE 97.91±0.07 69.74±0.21 97.76±0.05 88.66±0.21 98.89±0.02 93.11±0.13
AUPRC DivDis 83.85±0.41 75.09±0.38 89.62±0.12 80.92±0.06 93.34±0.10 81.48±0.30
FixMatch 80.92±1.12 70.57±0.32 87.14±0.58 81.51±0.31 92.38±0.15 82.78±0.57
D-BAT 76.64±0.49 54.87±0.99 84.70±0.51 70.08±0.69 90.84±0.28 73.45±0.90
AdvStyle 81.17±1.72 58.40±2.18 83.01±0.90 69.48±2.16 88.54±1.44 72.11±1.49
EoA 64.16±0.51 36.50±1.48 69.66±0.47 57.71±0.79 79.12±0.09 56.52±0.42
Mixup 73.03±1.67 60.84±4.31 80.36±0.88 72.88±1.67 86.88±0.14 74.99±0.15
NCDG 72.74±0.05 55.40±0.20 83.09±0.03 70.63±0.17 89.91±0.03 69.14±0.05
EMOE base 88.44±0.05 71.80±0.07 91.21±0.02 82.61±0.05 94.91±0.03 84.13±0.06
EMOE 89.52±0.05 71.41±0.11 91.59±0.03 83.06±0.10 95.38±0.01 84.77±0.02
AUROC DivDis 65.41±0.81 58.53±0.46 66.68±0.28 57.04±0.08 73.23±0.40 61.15±0.50
FixMatch 70.32±0.59 52.69±0.69 66.37±0.50 58.37±0.58 74.38±0.16 62.05±0.55
D-BAT 75.26±0.28 58.21±0.26 72.09±0.19 60.32±0.25 80.31±0.08 64.82±0.18
AdvStyle 75.97±0.39 58.86±0.25 70.78±0.35 59.62±0.30 78.36±0.23 64.14±0.27
EoA 64.91±0.34 52.71±0.44 59.27±0.20 54.63±0.24 62.99±0.16 55.83±0.18
Mixup 68.20±0.66 56.33±0.45 60.39±0.40 56.50±0.37 64.24±1.23 57.75±0.80
NCDG 73.70±0.09 56.48±0.06 70.17±0.04 59.77±0.02 77.78±0.09 63.14±0.06
EMOE base 73.70±0.07 56.48±0.06 70.17±0.04 59.77±0.02 77.78±0.09 64.89±0.06
EMOE 75.47±0.09 55.40±0.20 71.28±0.10 60.63±0.17 80.01±0.03 66.14±0.05
B.1 Full Experiment Results on CheMBL and Therapeutics Data Commons
In Tab. 7, we report the full results on hERG, A549_cells, cyp_2D6. and Ames.
Table 7: Full experiment results on ChEMBL (Gaulton et al. 2011) and Therapeutics Data Commons (Huang et al. 2021) datasets. We bold best scores based on the mean minus 1 standard deviation. Note that * refers to a semi-supervised method.
hERG A549_cells cyp_2D6 Ames
AUPRC DivDis 86.92±8.44 89.01±3.53 83.65±6.17 97.47±1.68
R<<0.1 FixMatch 84.77±1.81 92.93±3.00 92.71±1.98 96.26±1.04
D-BAT 88.55±1.68 98.57±0.16 95.71±0.89 99.07±0.23
AdvStyle 93.27±0.56 96.89±0.30 84.21±2.07 99.52±0.12
EoA 63.80±0.42 61.31±0.31 61.77±0.41 78.74±0.43
Mixup 82.80±1.56 95.04±0.25 87.39±3.09 91.02±1.05
NCDG 72.96±1.31 78.79±2.35 61.77±1.15 89.68±0.38
EMOE base 96.65±0.22 99.74±0.05 99.81±0.11 98.73±0.27
EMOE 98.10±0.34 99.76±0.03 99.48±0.17 99.66±0.20
AUPRC@ DivDis 85.65±2.37 89.45±1.16 79.98±1.10 95.80±0.34
R<<0.2 FixMatch 79.66±0.57 90.81±4.30 83.58±2.55 93.55±0.64
D-BAT 84.48±1.74 98.26±0.14 91.40±0.98 99.04±0.24
AdvStyle 88.21±0.77 97.77±0.28 84.83±0.93 99.05±0.17
EoA 63.80±0.42 61.31±0.31 61.77±0.41 78.74±0.43
Mixup 82.25±1.51 95.04±0.25 87.09±2.32 91.02±1.05
NCDG 70.25±1.03 78.22±2.13 61.48±0.98 86.29±0.45
EMOE base 94.44±0.17 98.22±0.0 95.51±0.19 97.84±0.20
EMOE 94.67±0.29 98.87±0.09 96.88±0.25 98.45±0.19
AUPRC@ DivDis 83.69±3.39 88.83±2.50 77.55±1.02 94.87±0.74
R<<0.3 FixMatch 76.16±1.04 89.12±5.07 77.70.58±2.05 92.72±0.51
D-BAT 82.44±1.59 97.37±0.24 87.65±0.58 98.61±0.24
AdvStyle 85.05±0.86 96.47±0.29 82.76±0.95 98.71±0.20
EoA 63.80±0.42 61.31±0.31 61.77±0.41 78.74±0.43
Mixup 81.51±1.35 94.95±0.25 84.53±2.89 90.88±1.00
NCDG 69.59±0.97 77.88±1.93 60.93±0.86 85.41±0.33
EMOE base 90.46±0.06 97.35±0.05 92.34±0.24 97.83±0.15
EMOE 90.88±0.34 97.96±0.1 93.06±0.27 98.18±0.15
AUPRC DivDis 67.70±0.25 76.45±0.63 65.76±0.45 82.37±0.27
FixMatch 45.16±3.11 51.31±1.28 32.65±2.16 72.54±1.15
D-BAT 54.60±1.59 67.04±0.53 47.42±0.85 70.44±0.76
AdvStyle 51.54±0.93 65.02±0.72 44.41±0.96 74.98±0.49
EoA 43.30±0.51 44.95±0.24 37.37±0.95 59.43±0.18
Mixup 42.42±0.85 50.52±0.50 27.79±1.49 60.94±0.80
NCDG 42.69±0.24 49.13±0.79 31.08±0.16 67.79±0.93
EMOE base 72.19±0.07 84.10±0.01 72.93±0.13 87.43±0.02
EMOE 73.26±0.08 84.60±0.08 73.59±0.15 88.50±0.10
AUROC DivDis 71.19±0.57 72.20±0.41 64.50±0.77 77.05±0.47
FixMatch 68.41±0.84 61.16±1.31 80.85±0.40 70.32±0.59
D-BAT 76.58±0.45 78.16±0.23 67.54±0.47 83.82±0.15
AdvStyle 75.84±0.46 76.13±0.28 65.51±0.62 85.56±0.71
EoA 68.02±0.34 68.33±0.24 60.50±0.48 74.77±0.25
Mixup 73.96±0.26 76.57±0.42 67.53±0.90 78.43±0.49
NCDG 70.42±0.13 72.50±0.79 62.01±0.16 73.42±0.93
EMOE base 74.74±0.06 79.17±0.02 70.33±0.07 81.87±0.04
EMOE 75.97±0.08 79.53±0.06 70.33±0.13 83.78±0.12
B.2 Full Experiment Results on DrugOOD
In Tab. 6, we report the full results on core ec50, refined ec 50, and core ic50 from DrugOOD (Ji et al. 2023).
Table 8: Full experiment results on Tableshift (Gardner, Popovic, and Schmidt 2023) datasets. We bold best scores based on the mean minus 1 standard error. Note that * refers to a semi-supervised method that uses additional unlabeled OOD data during training.
Childhood FICO Hospital Sepsis
Lead HELOC Readmission
AUPRC@ DivDis 91.67±3.03 84.88±3.80 88.99±3.36 59.50±1.55
R<<0.1 FixMatch 86.68±3.32 83.99±2.27 73.97±2.64 17.84±1.87
D-BAT 52.67±0.00 92.97±0.59 83.73±0.12 81.99±0.33
AdvStyle 63.63±0.00 90.10±1.03 77.29±0.74 60.79±0.72
EoA 75.73±0.83 63.42±2.13 52.03±1.78 43.76±1.14
Mixup 50.00±0.00 92.79±0.03 85.71±0.35 68.21±1.28
NCDG 83.30±0.20 91.70±0.33 70.42±0.17 72.40±0.04
EMOE base 98.69±0.03 92.18±0.58 51.31±0.16 82.85±1.53
EMOE 99.72±0.10 93.93±0.87 66.65±0.17 80.74±1.77
AUPRC@ DivDis 89.77±2.51 85.50±2.65 83.22±2.57 60.66±1.57
R<<0.2 FixMatch 81.75±3.69 77.76±1.94 67.84±1.95 17.28±0.58
D-BAT 62.82±0.00 91.20±0.24 78.84±0.12 75.37±0.38
AdvStyle 64.96±0.01 88.71±0.65 72.91±0.58 59.83±0.63
EoA 77.43±0.97 59.53±2.24 51.83±1.66 41.10±1.56
Mixup 50.00±0.00 91.16±2.10 69.19±3.95 66.09±1.15
NCDG 82.49±0.18 90.93±0.31 68.88±0.15 70.22±0.03
EMOE base 97.39±0.06 90.07±0.32 58.30±0.09 78.62±1.57
EMOE 97.92±0.20 91.72±0.62 67.57±0.12 76.95±1.45
AUPRC@ DivDis 87.90±2.20 85.77±2.73 79.82±2±42 61.55±1.55
R<<0.3 FixMatch 81.61±3.09 76.48±4.13 64.47±1.70 18.45±0.46
D-BAT 68.11±0.00 90.11±0.13 75.26±0.10 70.92±0.44
AdvStyle 68.67±0.08 87.33±1.21 70.27±0.49 59.02±0.57
EoA 79.88±1.03 56.52±2.34 51.34±1.27 40.09±1.05
Mixup 50.00±0.00 90.01±0.02 47.18±3.54 64.41±0.95
NCDG 82.10±0.15 88.76±0.29 67.52±0.13 69.48±0.03
EMOE base 96.08±0.08 89.73±0.23 60.86±0.04 75.84±1.55
EMOE 96.58±0.22 91.10±0.41 67.36±0.06 74.02±1.20
AUPRC DivDis 76.33±0.86 82.47±1.05 65.95±2.15 58.85±1.21
FixMatch 75.39±0.58 72.22±4.46 56.54±1.21 18.24±0.66
D-BAT 71.85±0.01 80.91±0.17 63.29±0.08 58.17±0.21
AdvStyle 48.37±2.77 79.63±1.65 38.13±3.80 54.34±0.27
EoA 49.48±0.19 59.53±2.24 29.45±4.95 11.21±2.21
Mixup 50.00±0.00 80.95±0.63 14.15±1.06 56.80±0.40
NCDG 73.08±0.14 79.05±0.21 58.95±0.10 61.96±0.02
EMOE base 86.39±0.19 83.80±0.07 62.83±0.03 64.01±0.94
EMOE 86.70±0.28 84.02±0.09 63.62±0.08 62.21±0.52
AUROC DivDis 77.74±1.90 83.55±1.06 65.28±0.59 62.33±0.61
FixMatch 79.87±0.29 74.29±1.69 55.78±0.98 49.50±1.01
D-BAT 79.13±0.02 76.13±0.03 63.22±0.03 57.95±0.04
AdvStyle 74.45±0.04 77.23±1.69 61.32±0.34 55.31±0.34
EoA 72.62±0.32 54.67±2.55 51.65±1.70 49.14±2.60
Mixup 50.00±0.00 78.74±0.17 63.37±0.25 56.82±0.26
NCDG 76.91±0.13 75.09±0.79 58.67±0.16 63.41±0.93
EMOE base 84.88±0.19 83.53±0.03 63.18±0.03 62.53±0.82
EMOE 87.95±0.25 83.11±0.04 63.65±0.07 61.42±0.35
B.3 Diversity of Predictions
In drug discovery applications, models should predict on structurally diverse compounds. To assess the diversity of model behavior in high confidence out-of-distribution (OOD) predictions, we examine the average variance of fingerprint features for instances with predicted confidence greater than 0.9 on the 3 ChEMBL datasets (var@p>0.9). Higher variance reflects greater heterogeneity among the selected molecules. We observe the following var@p>0.9 on ChEMBL datasets: EMOE (0.391), D-BAT (0.337), EoA (0.301), AdvStyle (0.349), Mixup (0.370), and NCDG (0.239). That is, EMOE is assigning confident predictions to structurally diverse compounds rather than overfitting to a narrow subset of the chemical space.
B.4 Full Experiment Results on Tableshift
In Table 8, we report the full results on full results on Tableshift (Gardner, Popovic, and Schmidt 2023) datasets.
B.5 Full Experiment Results on PACS Dataset
The PACS dataset includes images from four distinct domains: photo, art painting, cartoon, and sketch. Specifically, we focus on the animal classes (dog, elephant, giraffe, and horse) to create 2 challenging binary classification tasks. Models were trained on the ‘photo’, ‘art’, and ‘cartoon’ domains; they were tested on the unseen fourth domain, ‘sketch’, to assess generalization performance. Results on the PACS are shown in In Table 9, we report the full results on images from PACS (Li et al. 2017).
Table 9: Experiment results on PACS dataset. We bold best scores based on the mean minus 1 standard deviation.
PACS dog-elephant PACS giraffe-horse
AUPRC D-BAT 58.33±7.50 82.15±3.10
$\mathrm{@R}\!\!\leq\!\!.1$ AdvStyle 54.52±7.86 88.73±7.30
EoA 44.84±6.95 69.41±3.66
Mixup 65.20±7.80 84.07±4.53
EMOE base 66.58±1.27 83.23±0.71
EMOE 64.00±1.73 85.90±2.92
AUPRC@ D-BAT 58.35±7.27 80.20±2.58
$\mathrm{@R}\!\!\leq\!\!.2$ AdvStyle 55.83±5.64 84.77±6.41
EoA 45.16±6.39 68.53±4.25
Mixup 56.46±6.74 81.51±5.45
EMOE base 60.36±0.39 82.86±0.36
EMOE 61.94±1.82 84.10±1.44
AUPRC@ D-BAT 57.88±5.46 78.21±2.54
$\mathrm{@R}\!\!\leq\!\!.3$ AdvStyle 56.25±4.83 81.76±6.03
EoA 45.27±6.21 67.57±3.79
Mixup 56.33±6.51 79.40±4.96
EMOE base 58.62±0.07 82.69±0.43
EMOE 60.80±1.44 83.00±1.35
AUPRC D-BAT 54.27±2.78 66.10±1.74
AdvStyle 53.52±2.00 67.94±3.99
EoA 45.42±5.95 68.40±4.41
Mixup 54.05±1.80 67.99±2.66
EMOE base 54.47±0.01 73.10±0.64
EMOE 56.64±1.47 72.76±1.25
AUROC D-BAT 56.44±2.66 63.96±2.38
AdvStyle 56.24±1.28 65.88±3.53
EoA 49.82±0.92 54.93±3.95
Mixup 57.39±1.16 64.12±2.78
EMOE base 56.94±0.15 72.86±0.61
EMOE 59.99±2.12 72.38±1.18
B.6 Base Experts Ablations
Table 10: Mean AUPRC on ChEMBL using D-BAT as EMOE Base.
hERG A549_cells cyp_2D6
AUPRC EMOE base D-BAT 86.29 98.37 96.66
@R<<0.1 EMOE 87.05 98.92 98.51
AUPRC EMOE base D-BAT 84.86 97.45 92.56
@R<<0.2 EMOE 85.66 99.05 95.33
AUPRC EMOE base D-BAT 84.12 96.97 89.06
@R<<0.3 EMOE 84.13 98.01 91.40
AUPRC EMOE base D-BAT 72.21 83.79 73.26
EMOE 72.73 84.80 73.01
In this section, We ablate the kind of base experts as EMOE NN relies on the base experts to provide pseudo-labels during training. reports the mean AUPRC@R<<0.2 across ChEMBL datasets when using XGBoost, Random Forest (with 100 estimators), and Decision Tree as base experts. Performance was comparable between XGBoost and Random Forest, indicating robustness to the choice of strong ensemble models. In contrast, using a weaker base model such as a single Decision Tree led to a performance drop. Nevertheless, EMOE consistently outperformed its respective base experts, demonstrating its ability to enhance predictions regardless of base model strength. In addition to the simple models, we conducted ablation on base experts with the most competitive baseline model D-BAT(Pagliardini et al. 2023). We trained 64 D-BAT models each with 5 models in them and apply the same EMOE loss and NN architecture training using this set of base models. Results shows EMOE is able to improve upon complex OOD generalization model like D-BAT (see Tab.10).
B.7 Per-head Matching Ablation Details
In this section we provide details on the per-head matching ablations where we ablate the matching loss scheme on base experts and explore a mean-only matching approach on expanded points as an alternative. First, we consider utilizing a single-headed (SH) MLP, $f(x)$ ($512\rightarrow 512\rightarrow 1$), which is trained via a mean matching loss $\mathcal{L}_{\mathrm{MM}}(f,\{g_{j}\}_{j=1}^{K};\mathcal{S})\equiv\frac{1}{|\mathcal{S}|}\sum_{x\in S}\ell(f(x),\frac{1}{K}\sum_{j=1}^{K}g_{j}(x))$, rather than the per-expert matching loss, $\mathcal{L}_{\mathrm{match}}$ (4) We also explored the effect of training our multi-headed (MH) architecture ($512\rightarrow 512\rightarrow 1024$) using only mean-matching (without per-head matching), $\mathcal{L}^{\prime}_{\mathrm{MM}}(\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{S})\equiv\frac{1}{|\mathcal{S}|}\sum_{x\in S}\ell(\frac{1}{K}\sum_{j=1}^{K}\sigma(h_{j}(x)),\frac{1}{K}\sum_{j=1}^{K}g_{j}(x)).$
B.8 Bottleneck Ablations Details
Table 11: ChEMBL datasets’ mean AUPRC ablating hidden layer and output sizes.
@R<<.2 @R<<1
Full 97.12 77.28
Full (ML) 96.94 77.25
Full (ERM) 92.65 72.61
Tiny 96.29 76.63
Tiny (ML) 96.33 76.17
Tiny (ERM) 88.28 70.87
In this section, we provide the details and results for the bottleneck ablations. Results are shown in Tab. 11, here ‘Full’ denotes our original $2\times 512$ hidden layer architecture, where ‘tiny’ denotes a $2\times 32$ hidden layer architecture (a $\times 16$ decrease in parameters). Moreover, we also consider mean logits ‘ML,’ a final averaging over the multi-head logits that produces a single output unit (i.e., averaging the output weights/bias after training to construct the mean logits network). The performance gap is marginal between the ‘Full’ and ‘Tiny’ model (a $0.85\%$ difference) when using our proposed loss. In contrast, when using empirical risk minimization, we see a $4.87$ times bigger drop in performance between ‘Full’ and ‘Tiny’ models. This suggests that the bottlenecking properties of our method are key to EMOE’s performance. Moreover, the results show promise for EMOE in resource-constrained settings (such as in IoT applications).
Appendix C Additional Figures
Predicted probabilities from EMOE network and base experts. We highlight example instances where the base experts initially makes incorrect predictions but are corrected when we average the predicted probabilities from EMOE network and EMOE base.
Appendix D Limitations
While EMOE consistently improves over its base experts across diverse model types (e.g., XGBoost, Random Forest, Decision Tree, D-BAT), its effectiveness can be influenced by the quality of the base models and the learned latent space. However, our results show that EMOE remains robust even with simpler base models and standard latent representations, suggesting room for further gains with more sophisticated choices. To maintain broad applicability, our experiments were constrained to real-valued vector data with general augmentations. In domain-specific applications, incorporating modality-aware augmentations could further enhance performance. Future work may explore this direction to extend EMOE’s effectiveness.